### v1

# A generalization of Frenkel’s formula

Abstract.
We generalize Frenkel’s integral formula for traces of operators to operators. The resulting formula holds for bounded self-adjoint positive operators and $p$-Schatten class of compact positive operators.
Key words and phrases:
Frenkel’s formula, divergence, operator divergence, $p$-Schatten operators.
2010 Mathematics Subject Classification:
81P47,94A40
1. Introduction
Denote: $\mathbb{C}^{n\times n}\supset\mathrm{H}_{n}\supset\mathrm{H}_{n,+}\supset\mathrm{H}_{n,++}$ the complex space of $n\times n$ matrices, the real space of $n\times n$ Hermitian matrices, the cone of positive semidefinite matrices, and the open set of positive definite matrices. Let $\mathrm{H}_{n,+,1}\supset\mathrm{H}_{n,++,1}$ be the convex set and the open subset of positive density matrices of order $n$ respetively. For $A,B\in\mathrm{H}_{n}$ we denote $\lambda_{1}(A)\geqslant\ldots\geqslant\lambda_{n}(A)$ the eigenvalues of $A$, and $A>B(A\geqslant B)$ if $A-B\in\mathrm{H}_{n,++}(A-B\in\mathrm{H}_{n,+})$. Here, for any $T\in\mathrm{H}_{n}$ we denote by $\{T>0\}$ the projection on the positive spectrum of $T$, and
(1) $T_{+}=T\{T>0\},\quad T_{-}=T\{-T>0\},\quad T=T_{+}-T_{-}.$
The von Neumann entropy of $\rho\in\mathrm{H}_{n,+,1}$ is $\operatorname{Tr}\rho\log\rho$, and it is equal to the Shannon entropy of the eigenvalue set of $\rho$. The Umegaki relative entropy of $\rho$ with respect to $\sigma$, (the divergence of $\rho$ with respect to $\sigma$), is $\operatorname{Tr}\rho(\log\rho-\log\sigma)$ [15]. It is convenient to define the divergence of $A,B\in\mathrm{H}_{n,+}$ as [4, eq. (3)]:
(2) $D(A\|B):=\operatorname{Tr}\big(A(\log A-\log B))-A+B\big).$
Furthermore, for $B>0$ let
(3) $D\log[B](A):=\lim_{t\to 0}\frac{\log(B+tA)-\log B}{t}.$
For numerical computations we have the following well known observations: the von Neumann entropy of $A$ can be computed by finding the eigenvalues of $A$; $D(A\|B)$ can be found by using the spectral decompositons of $A$ and $B$. A less known observation is that $D\log[B](A)$ can be found by using the spectral decomposition of $B$ as follows: Find a unitary matrix $U$ such that $U^{*}BU=\operatorname{diag}(b_{1},\ldots,b_{n})$ and let $U^{*}AU=[c_{ij}]$ be Hermitian matrices of order $n$. Then [2, Theorem V.3.3] or [7, eq. (5)] state:
(4) $D\log[B](A)=U([c_{ij}\frac{\log b_{i}-\log b_{j}}{b_{i}-b_{j}}]U^{*},\quad\frac{\log b-\log a=b}{b-b}=\frac{1}{b}\textrm{ for }b>0.$
Observe that
(5) $\operatorname{Tr}BD\log[B](A)=\operatorname{Tr}A,\quad D\log[B](B)=I_{n},$
where $I_{n}\in\mathrm{H}_{n}$ is the identity matrix. In many theoretical applications physicists prefer integral formulas for the above and similar functions. A new notable formula that attracted recent activity in quantum information theory (QIT) is Frenkel’s integral formula for $D(A\|B)$ [5]:
(6) $D(A\|B)=\int_{-\infty}^{\infty}\frac{dt}{|t|(t-1)^{2}}\operatorname{Tr}\big((1-t)A+tB\big)_{-}$
See for example [3, 4, 10, 12].
The aim of this note is to give the following generalization of (6) to operators:
(7) $A(\log A-\log B)-BD\log[B](A)+B=\int_{-\infty}^{\infty}\frac{dt}{|t|(t-1)^{2}}\big((1-t)A+tB\big)_{-},\quad A\geqslant 0,B>0.$
We call the left hand side of this formula the divergence operator:
(8) $\Delta(A\|B):=A(\log A-\log B)-BD\log[B](A)+B.$
Let
(9) $O_{\gamma}(A\|B)=(A-\gamma B)_{+}.$
We show the following equivalent formula analogous to [12, Proposition 4.2] and [4, eq. (10)]:
(10) $\Delta(A\|B)=\int_{1}^{\infty}\big(\gamma^{-1}O_{\gamma}(A\|B)+\gamma^{-2}O_{\gamma}(B\|A)\big)d\gamma,\quad A\geqslant 0,B>0.$
It is easy to verify these equalities if $AB=BA$, by assuming that both $A$ and $B$ are diagonal matrices. The surprising fact that these equalites hold for any $A,B>0$. We suspect that there are other integral formulas in QIT where the trace can be replaced by the corresponding operators.
2. Proof of (7)
Corollary 1 yields that for $A,B\in\mathrm{H}_{n}$, the matrix $(A+tB)_{+}$, hence $(A+tB)_{-}=(A+tB)_{+}-A-tB$, is continuous in $t$ on the real line $\mathbb{R}$. Furthermore, there exists as finite number of points $Z\subset\mathbb{R}$, possibly an empty set, such that $(A+tB)_{+}$ is analytic on each connected interval of $\mathbb{R}\setminus Z$. Hence all the integrals in section 1 make sense, provided that they converge.
As in [10, 12, 4] we give an alternative form of the right hand side of (7).
(11) $\int_{-\infty}^{\infty}\frac{dt}{|t|(t-1)^{2}}\big((1-t)A+tB\big)_{-}=\int_{1}^{\infty}\big(\gamma^{-1}O_{\gamma}(A\|B)+\gamma^{-2}O_{\gamma}(B\|A)\big)d\gamma.$
First observe that if $A\geqslant 0,B\geqslant 0$ then $((1-t)A+tB)_{-}=0$ for $t\in[0,1]$. Hence,
$\int_{0}^{\infty}\frac{dt}{t(t-1)^{2}}\big((1-t)A+tB\big)_{-}=\int_{1}^{\infty}\frac{dt}{t(t-1)^{2}}\big((1-t)A+tB\big)_{-}.$
Introduce a new variable $\gamma=\frac{t}{t-1}$ on the interval $(1,\infty)$, and recall that $(-C)_{-}=C_{+}$ to deduce
$\frac{dt}{t(t-1)^{2}}\big((1-t)A+tB\big)_{-}=\frac{dt}{t(t-1)^{2}}\left((t-1)(-1)\big(A-\frac{t}{t-1)}B\big)\right)_{-}=$
$\frac{dt}{t(t-1)}\big(A-\gamma B\big)_{+}=-\frac{d\gamma}{\gamma}\big(A-\gamma B\big)_{+}.$
Hence,
$\int_{0}^{\infty}\frac{dt}{t(t-1)^{2}}\big((1-t)A+tB\big)_{-}=\int_{1}^{\infty}\frac{O_{\gamma}(A\|B)}{\gamma}d\gamma.$
On the interval $(-\infty,0)$ we introduce a new variable $\gamma=-\frac{1-t}{t}$. Observe
$\big((1-t)A+tB\big)_{-}=-t\big(B-\gamma A)_{+}\textrm{ for }t<0\Rightarrow$
$\int_{-\infty}^{0}-\frac{dt}{t(1-t)^{2}}\big((1-t)A+tB\big)_{-}=\int_{1}^{\infty}\frac{d\gamma}{\gamma^{2}}O_{\gamma}(B\|A).$
This establishes (11).
To prove (7) we follow the proof of Frenkel’s integral formula in [12, Section 4.1]. Recall [12, eq. (77)] and [3, Theorem B.1] for $A,B>0$:
(12) $\log A-\log B=\int_{1}^{\infty}\big(\{A-\gamma B>0\}-\{B-\gamma A>0\}\big)\frac{1}{\gamma}d\gamma,$
$D\log[A](B)=\int_{0}^{\infty}\{B-\gamma A>0\}d\gamma.$
Hence
(13) $A(\log A-\log B)=\int_{1}^{\infty}\big(A\{A-\gamma B>0\}-A\{B-\gamma A>0\}\big)\frac{1}{\gamma}d\gamma=u+v-w,$
$u=\int_{1}^{\infty}\big(\gamma^{-1}(A-\gamma B)\{A-\gamma B>0\}+\gamma^{-2}(B-\gamma A)\{B-\gamma A>0\}\big)d\gamma,$
$v=\int_{1}^{\infty}B\{A-\gamma B>0\}d\gamma,\quad w=\int_{1}^{\infty}\gamma^{-2}B\{B-\gamma A>0\}d\gamma.$
Clearly, $u$ is equal to the right hand side of (11). Use the second equality of (12) to deduce that
$v=BD\log[B](A)-\int_{0}^{1}B\{A-\gamma B>0\}d\gamma.$
We now evaluate $w$. Observe that on the interval $(0,1)$ one has the equality $\{-A+\alpha B>0\}=I_{n}-\{A-\alpha B>0\}$ except a finite number of points. (Here, $I_{n}$ is the identity matrix of order $n$.) Hence, by changing to the variable $\alpha=\gamma^{-1}$ for $\gamma\in(1,\infty)$ we obtain
$w=\int_{0}^{1}B\{\alpha B-A>0\}d\alpha=\int_{0}^{1}B(I_{n}-\{A-\alpha B>0\}d\alpha=B-\int_{0}^{1}B\{A-\alpha B>0\}d\alpha.$
Combine all the above equalities to deduce (7) for $A,B>0$. Fix $B>0$ and let $A_{k}>0,k\in N$ and let $\lim_{k\to\infty}A_{k}=A\geqslant 0$. This shows that (7) holds $A\geqslant 0$ and $B>0$. One can generalize further as in [3, Theorem C.1] to matrices to
(14) $0\leqslant A\leqslant\tau B\textrm{ for some }\tau\geqslant 1.$
For completeness we bring the following lemma
Lemma 1.
Let $A\in\mathrm{H}_{n,+},B\in\mathrm{H}_{n}$. Then
  1. (1)
    The integral $f(t):=\int_{1}^{t}\gamma^{-2}O_{\gamma}(B\|A)d\gamma$ converges to $F\in\mathrm{H}_{n,+}$ as $t\to\infty$.
  2. (2)
    Assume that (14) holds. Then
    $\int_{1}^{\infty}\gamma^{-1}O_{\gamma}(A\|B)d\gamma=\int_{1}^{\tau}\gamma^{-1}O_{\gamma}(A\|B)d\gamma.$
  3. (3)
    Assume that $A,B\geqslant 0$ and (14) does not hold. Then, $\operatorname{supp}A\not\subseteq\operatorname{supp}B$. That is, there exists $\mathbf{x}\in\mathbb{C}^{n},$ such that $B\mathbf{x}=\mathbf{0}$ and $\mathbf{x}^{*}A\mathbf{x}>0,\mathbf{x}^{*}B\mathbf{x}=0$. Hence,
    (15) $\lim_{t\to\infty}\|\int_{1}^{t}\big(\gamma^{-1}O_{\gamma}(A\|B)+\gamma^{-2}O(B\|A)\big)d\gamma\|=\infty.$
Proof.
(1) Corollary 1 yields that $O_{\gamma}(A\|B),O_{\gamma}(B\|A)$ are continuous on $\mathbb{R}$. Hence $f(t)$ is well defined. Observe next that for $A\geqslant 0,\gamma\geqslant 0$:
(16) $0\leqslant O_{\gamma}(B\|A)=(B-\gamma A)_{+}=\{B-\gamma A>0\}(B-\gamma A)\{B-\gamma A>0\}=$
$\{B-\gamma A>0\}B\{B-\gamma A>0\}-\gamma\{B-\gamma A>0\}A\{B-\gamma A>0\}\leqslant$
$\{B-\gamma A>0\}B\{B-\gamma A>0\}\leqslant\|B\|I_{n}\Rightarrow\|O_{\gamma}(B\|A)\|\leqslant\|B\|$
Hence,
$1\leqslant s\leqslant t\Rightarrow\|f(t)-f(s)\|\leqslant\|B\|\int_{s}^{t}\gamma^{-2}d\gamma\leqslant\|B\|\int_{s}^{\infty}\gamma^{-2}d\gamma=\frac{\|B\|}{s}.$
Therefore, $f(m),m\in\mathbb{N}$ is a Cauchy sequence that converges to $F\in\mathrm{H}_{n,+}$. Furthermore, the above inequality proves (1).
(2) Assume (14) holds. For $t\geqslant\tau$ we have
$A-\tau B=A-\tau B-(t-\tau)B\leqslant 0\Rightarrow(A-\tau B)_{+}=0\Rightarrow\int_{1}^{\infty}\gamma^{-1}O_{\gamma}(A\|B)d\gamma=\int_{1}^{\tau}\gamma^{-1}O_{\gamma}(A\|B)d\gamma.$
(3) Assume that $A,B\geqslant 0$ and (14) does not hold. Hence $B$ is singular. Let $\mathbf{V}\subset\mathbb{C}^{n}$ be the eigenspace of $B$ corresponding to 0-eigenvalue. Suppose that the sesquilinear form $\mathbf{x}^{*}A\mathbf{x}$ vanishes on $\mathbf{V}$. As $A\geqslant 0$ it follows that $A\mathbf{V}=\{\mathbf{0}\}$. By restricting $A$ and $B$ to the orthogonal complement $\mathbf{V}^{\perp}$ of $\mathbf{V}$ we deduce that (14) holds contrary to ours assumption. Hence, there exists $\mathbf{x}\in\mathbf{V}$ such that $B\mathbf{x}=0$ and $\mathbf{x}^{*}A\mathbf{x}>0$. Therefore,
$0<\mathbf{x}^{*}A\mathbf{x}=\mathbf{x}^{*}(A-\gamma B)\mathbf{x}\leqslant\mathbf{x}^{*}(A-\gamma B)_{+}\mathbf{x}\Rightarrow\lim_{t\to\infty}\int_{1}^{t}\gamma^{-1}\mathbf{x}^{*}O(A\|B)_{\gamma}\mathbf{x}d\gamma=\infty.$
Theorem 1.
Assume that $A,B\in\mathrm{H}_{n,+}$. Then the following dichotomy holds:
  1. (1)
    Assume that there exists $\tau\geqslant 1$ such that $A\leqslant\tau B$. Then $\Delta(A\|B)\in\mathrm{H}_{n,+}$, and equality (10) holds,where
    (17) $\Delta(A\|B)=\Delta(A_{1}\|B_{1})\oplus 0,\quad A_{1}=A|B\mathbb{C}^{n},B_{1}=B|B\mathbb{C}^{n}.$
  2. (2)
    Assume that there is no $\tau\geqslant 1$ such that $A\leqslant\tau B$. Then the right side of (10) converges to an unbounded nonnegative operator. More precisely, there exists $\mathbf{x}\in\mathbb{C}^{n}$ such that $B\mathbf{x}=0$ and $\mathbf{x}^{*}A\mathbf{x}>0$. Hence, (15) holds. In particular, in equality (6) both sides are equal to $\infty$.
Proof.
Lemma 1 yields that it is enough to consider the case where $B$ is singular and $A\leqslant\tau B$ for some $\tau\geqslant 1$. Let $\mathbf{V}\subset\mathbb{C}^{n}$ be the kernel of $B$. Observe that $A\mathbf{V}=\{\mathbf{0}\}$. Thus, we can decompose $A=A_{1}\oplus 0,B=B_{1}\oplus 0$, where $A_{1},B_{1}$ are the restrictions of $A$ and $B$ to $B\mathbb{C}^{n}$. Then $\Delta(A\|B)$ is defined by (17). ∎
3. An infinite dimensional case
Let $\mathcal{H}$ be an infinite dimensional separable Hilbert space with an inner product $\langle\cdot,\cdot\rangle$, and the norm $\|\mathbf{x}\|=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle}$. Denote by $\mathcal{B}\supset\mathcal{S}\supset\mathcal{S}_{+}\supset\mathcal{S}_{++}$ the sets of bounded, self-adjoint, nonnegative definite, and positive definite operators on $\mathcal{H}$. For $T\in\mathcal{B}$ let $\|T\|=\sup_{\mathbf{x},\|\mathbf{x}\|=1}\|T\mathbf{x}\|$. Assume that $T\in\mathcal{S}$. As in the finite dimenisional case denote by $\{T>0\}$ the projection on the positive spectrum of $T$. See [13]. Then the spectral decomposition of $T$ yields the equality (1). For $T\in\mathcal{B}$ let $|T|=\sqrt{T^{*}T}$. Then
(18) $T=T_{+}-T_{-},\quad|T|=T_{+}+T_{-},$
$T_{+}=\frac{1}{2}(|T|+T),\quad T_{-}=\frac{1}{2}(|T|-T),$
$\|T_{+}\|=\sup_{\|x\|\leqslant 1}\langle\mathbf{x},T\mathbf{x}\rangle,\quad\|T_{-}\|=\sup_{\|x\|\leqslant 1}-\langle\mathbf{x},T\mathbf{x}\rangle,\quad\|T\|=\sup_{\mathbf{x}\in\mathcal{H},\|\mathbf{x}\|=1}|\langle\mathbf{x},T\mathbf{x}\rangle|.$
Denote by $\mathcal{K}_{p}\supset\mathcal{S}_{p}\supset\mathcal{S}_{p,+}\supset\mathcal{S}_{p,++}$ the $p$-Schatten space of compact operators, the subspace of self-adjoint, nonnegative definite, and positive definite operators with the norm $\|T\|_{p}=\big(\operatorname{Tr}|T|^{p}\big)^{1/p},p\in[1,\infty]$.
It is stated in [10, Section 2.2] that (6) holds for $A,B\in\mathcal{S}_{1,+}$ as follows. Assume first that $\operatorname{supp}A\not\subseteq\operatorname{supp}B$. Then as in Theorem 1 both sides of (1) are $\infty$. Assume second that $\operatorname{supp}A\subseteq\operatorname{supp}B$. As in Theorem 1 we can restrict $A$ and $B$ to $\operatorname{range}B$. Thus, it is enough to discuss the case where $B>0$. Thus, we can summarize [10, Theorem 1] in this form:
Theorem 2.
Assume that $A\in\mathcal{S}_{1,+},B\in\mathcal{S}_{1,++}$. Let
(19) $e_{1}:=\int_{1}^{\infty}\gamma^{-1}\|O_{\gamma}(A\|B)\|_{1}d\gamma+\int_{1}^{\infty}\gamma^{-2}\|O_{\gamma}(B\|A)\|_{1}d\gamma.$
Then the following dichotomy holds: Either $e_{1}<\infty$, in which case both sides of (6) are finite and equal, or $e_{1}=\infty$ and both sides of (6) are $\infty$.
We now give the following versions of Theorem 1:
Theorem 3.
[…]
(20) $A_{n}=P_{n}AP_{n},\quad B_{n}=P_{n}BP_{n},$
$C_{n}|\mathbf{V}_{n}=A_{n}(\log A_{n}-\log B_{n})|\mathbf{V}_{n},D_{n}=B_{n}D[B_{n}](A_{n})|\mathbf{V}_{n},\quad C_{n}|\mathbf{V}_{n}^{\perp}=D_{n}|\mathbf{V}_{n}^{\perp}=0.$
  1. (1)
    Let
    (21) $e:=\int_{1}^{\infty}\gamma^{-1}\|O_{\gamma}(A\|B)\|d\gamma+\int_{1}^{\infty}\gamma^{-2}\|O_{\gamma}(B\|A)\|d\gamma,\quad p\in[1,\infty].$
    Assume that $e<\infty$. Then (10) holds. The operator divergence $\Delta(A\|B)$ is defined as follows. The sequences $\{B_{n}\},\{C_{n}-D_{n}\}$ of finite dimensional operators converge in the strong topology to operators $B\in\mathcal{S}_{++},C-D\in\mathcal{S}$ respectively, where
    (22) $\Delta(A\|B):=C-D+B.$
    If there exists $\tau\geqslant 1$ such that $A\leqslant\tau B$ then $e<\infty$.
  2. (2)
    Fix $p\in[1,\infty]$, and assume that $A\in\mathcal{S}_{p,+},B\in\mathcal{S}_{p,++}$ Let
    (23) $e_{p}:=\int_{1}^{\infty}\gamma^{-1}\|O_{\gamma}(A\|B)\|_{p}d\gamma+\int_{1}^{\infty}\gamma^{-2}\|O_{\gamma}(B\|A)\|_{p}d\gamma,\quad p\in[1,\infty].$
    Assume that $e_{p}<\infty$. Then (10) holds. The left hand side of (10) is defined as follows. The sequences $\{B_{n}\},\{C_{n}-D_{n}\}$ of finite dimensional operators converge in the norm $\|\cdot\|_{p}$ to operators $B,C-D\in\mathcal{S}_{p}$, respectively, where (22) holds. If there exists $\tau\geqslant 1$ such that $A\leqslant\tau B$ then $e_{p}<\infty$ in $1,\infty]$.
Proof.
(1) Kato’s inequality (34) yields that $O_{\gamma}(A\|B),O_{\gamma}(B\|A)$ are continuous in $\gamma\in\mathbb{R}$. As $e_{\infty}<\infty$ it follows that $G(t)=\int_{1}^{t}\big(\gamma^{-1}O_{\gamma}(A\|B)+\gamma^{-2}O_{\gamma}(B\|A)\big)d\gamma$ converges to $G\in\mathcal{S}_{+}$ as $t\to\infty$ in the operator norm $\|\cdot\|$. Let $A_{n}=P_{n}AP_{n}\mathrm{H}_{n,+},B_{n}=P_{n}BP_{n}\in\mathrm{H}_{n,++}$, and
(24) $G_{n}(t)=\int_{1}^{t}\big(\gamma^{-1}O_{\gamma}(A_{n}\|B_{n})+\gamma^{-2}O_{\gamma}(B_{n}\|A_{n})\big)d\gamma.$
The inequality (35) yields
$\|O_{\gamma}(A_{n}\|B_{n})\|\leqslant\|O_{\gamma}(A\|B)\|,\quad\|O_{\gamma}(A_{n}\|B_{n})\|\leqslant\|O_{\gamma}(A\|B)\|\Rightarrow$
$\int_{1}^{\infty}\big(\gamma^{-1}O_{\gamma}(A_{n}\|B_{n})+\gamma^{-2}O_{\gamma}(B_{n}\|A_{n})\big)d\gamma\leqslant e.$
Hence, $G_{n}(t)=\int_{1}^{t}\big(\gamma^{-1}O_{\gamma}(A_{n}\|B_{n})+\gamma^{-2}O_{\gamma}(B_{n}\|A_{n})\big)d\gamma$ converges in the norm $\|\cdot\|$ to $G_{n}=\int_{1}^{\infty}\big(\gamma^{-1}O_{\gamma}(A_{n}\|B_{n})+\gamma^{-2}O_{\gamma}(B_{n}\|A_{n})\big)d\gamma$.
We claim that $G_{n}$ converges in the strong topology to $G_{n}$. Fix $\mathbf{y}\in\mathcal{H},\|\mathbf{y}\|=1$. Observe that
$|(G-G_{n})\mathbf{y}\|\leqslant\int_{1}^{\infty}\big(\gamma^{-1}\|(O_{\gamma}(A\|B)-O_{\gamma}(A_{n}\|B_{n}))\mathbf{y}\|+\gamma^{-2}\|(O_{\gamma}(B\|A)-O_{\gamma}(B_{n}\|A_{n}))\mathbf{y}\|\big)d\gamma\leqslant 2e$
For $\gamma\in[1,\infty)$ set
$f_{n}(\gamma)=\gamma^{-1}\|(O_{\gamma}(A\|B)-O_{\gamma}(A_{n}\|B_{n}))\mathbf{y}\|+\gamma^{-2}\|(O_{\gamma}(B\|A)-O_{\gamma}(B_{n}\|A_{n}))\mathbf{y}\|,\quad n\in\mathbb{N},$
$g(\gamma)=2\big(\gamma^{-1}\|\big(O_{\gamma}(A\|B)\mathbf{y}\|+\gamma^{-2}\|O_{\gamma}(B\|A)\mathbf{y}\|\big).$
Clearly, $f_{n}(\gamma)$, and $g(\gamma)$ are continuous funcrions, and $0\leqslant f_{n}(\gamma)\leqslant g(\gamma)$ for $\gamma\in[1,\infty]$ and $n\in\mathbb{N}$. The equality (35) yields that $\lim_{n\to\infty}f_{n}(\gamma)=0$. Lebesgue’s dominated convergence theorem yields that $G_{n},n\to\infty$ converges in the strong topology to $G$. As $B\in\mathcal{S}_{++}$ it follows that $B_{n}$ is positive definite on $\mathcal{H}_{n}=P_{n}\mathcal{H}$. Hence, part (1) of Theorem 1 holds for $A_{n}|\mathcal{H}_{n}$ and $B_{n}|\mathcal{H}_{n}$. As $G_{n},n\to\infty$ converges in the strong topology to $G$, we deduce that $C_{n}-D_{n}=B_{n}$ converges in the strong topology to $G$. The equality (37) yields that $A_{n}$ and $B_{n}$ converge in the strong topology to $A$ and $B$ respectively. Hence, $C_{n}-D_{n}$ converge in the strong topology to CDC-D
Assume now that $A\leqslant\tau B$ for some $\tau\geqslant 1$. Then
$e^{\prime}:=\int_{1}^{\infty}\gamma^{-1}\|O_{\gamma}(A\|B)\|d\gamma=\int_{1}^{\tau}\gamma^{-1}\|O_{\gamma}(A\|B)\|d\gamma<\infty.$
As in the proof of part (1) of Lemma 1 we deduce that $e^{\prime\prime}:=\int_{1}^{\infty}\gamma^{-2}O_{\gamma}(B\|A)d\gamma<\infty$.
(2) Fix $p\in[1,\infty]$. Let $L_{0},L_{1}\in\mathcal{S}_{p}$, and $L(\gamma):=L_{0}-\gamma L_{1}\in\mathcal{S}_{p}$ for $\gamma\in(-\infty,\infty)$. Then
$\|L(\gamma)\|_{p}=\big(\|L_{+}(\gamma)\|_{p}^{p}+\|L_{-}(\gamma)\|_{p}^{p}\big)^{1/p}\leqslant\|L_{0}\|_{p}+|\gamma|\|L_{1}\|_{p}.$
Recall Rellich’s theorem [9, Theorem 3.9]. The eigenvalues and the eigenvectors of the pencil $L(\gamma)$ are analytic on $\mathbb{R}$. Hence each nonzero analytic eigenvalue of $L(\gamma)$ has a countable number of zeros that accumulate to $\pm\infty$. Therefore $\|L_{+}(\gamma)\|_{p}$ is a continuous function on $\mathbb{R}$. (See Theorem 4 for the case where $L_{0},L_{1}$ have finite dimensional range.) Define
$f_{n,p}(\gamma)=\gamma^{-1}\|(O_{\gamma}(A\|B)-O_{\gamma}(A_{n}\|B_{n}))\|_{p}+\gamma^{-2}\|(O_{\gamma}(B\|A)-O_{\gamma}(B_{n}\|A_{n}))\|_{p},\quad n\in\mathbb{N},$
$h_{n,p}(\gamma)=\big(\gamma^{-1}\|\big(O_{\gamma}(A_{n}\|B_{n})\|_{p}+\gamma^{-2}\|O_{\gamma}(B_{n}\|A_{n})\|\big),$
$g_{p}(\gamma)=2\big(\gamma^{-1}\|\big(O_{\gamma}(A\|B)\|_{p}+\gamma^{-2}\|O_{\gamma}(B\|A)\|_{p}\big).$
Note that all the above functions are continuous on $\mathbb{R}$. The equalities (36) yield that the sequence $h_{n,p}(\gamma),n\in\mathbb{N}$ is an increasing sequence in $n\in\mathbb{N}$ for a fixed $\gamma$, which converges to $g_{p}(\gamma)/2$. Furthermore, $\lim_{n\to\infty}f_{n,p}(\gamma)=0$.
Assume that $e_{p}<\infty$. Then $G_{p}(t)=\int_{1}^{t}\big(\gamma^{-1}O_{\gamma}(A\|B)+\gamma^{-2}O_{\gamma}(B\|A)\big)d\gamma$ converges to $G_{p}\in\mathcal{S}_{p,+}$ as $t\to\infty$ in the norm $\|\cdot\|_{p}$. Let $G_{n,p}=\int_{1}^{\infty}\big(\gamma^{-1}O_{\gamma}(A\|B)+\gamma^{-2}O_{\gamma}(B\|A)\big)d\gamma\in\mathcal{S}_{p,+}$. Lebesgue’s dominated convergence theorem yield that $G_{n,p},n\to\infty$ converges in $\|\cdot\|_{p}$ to $G_{p}$. Hence, $C_{n}-D_{n}+B_{n}$ converge to $\Delta(A\|B)$. Equalities (36) imply that $\lim_{n\to\infty}\|B-B_{n}\|_{p}=0$. Hence, $C_{n}-D_{n}$ converges to $\Delta(A\|B)-B$.
Assume that $A\leqslant\tau B$. Then $e_{p}<\infty$ as in the proof of part (1). ∎
Acknowledgment
The author thanks Gilad Gour for pointing out the references [4, 5].
Appendix A Auxiliary results
A.1. A finite dimensional case
Definition 1.
A Hermitian pencils is defined:
(25) $H(z)=A+zB,\quad A,B\in\mathrm{H}_{n}.$
A pencil $H(z)$ is called nonsingular if $\det H(z)\not\equiv 0$, and $1$-indecomposable if $H(z)$ does not have a constant eigenvector: $\mathbf{x}=(a+bz)\mathbf{x},\mathbf{x}\neq 0. H(z)H(z)$ is called completely $1$-decomposable, if there exists a unitary $U\in\mathbb{C}^{n\times n}$ such that $H(z)=U\operatorname{diag}(a_{1}+zb_{1}),\ldots,a_{n}+zb_{n})U^{*}$.
Recall that $H(z)$ is completely $1$-decomposable if and only if $AB=BA$.
Lemma 2.
Assume that $H(z)$ is of the form (25).
  1. (1)
    The pencil $H(z)$ is either $1$-indecomposable, completely $1$-decomposable, or the there exists unitary $U\in\mathbb{C}^{n\times n}$ such
    (26) $H(z)=U\big(H_{1}(z)\oplus H_{2}(z)\big)U^{*},$
    $H_{2}(z)=\operatorname{diag}(a_{1}+zb_{1},\ldots,a_{k}+zb_{k})\big),k\in[n-2],\quad H_{1}(z))\in\mathbb{C}^{(n-k)\times(n-k)}[z],$
    where $H_{1}(z)$ is 1-indecomposable.
  2. (2)
    Assume that $H(z)$ is 1-indecomposable. Then either $H(z)$ is nonsingular, or there exists an invertible $P\in\mathbb{C}^{n\times n}$ such that
    (27) $H(z)=P\big(\oplus_{j=1}^{k+1}H_{j}(z)\big)P^{*},$
    $H_{j}(z)=\begin{bmatrix}0_{m_{j}\times m_{j}}&L_{m_{j}}(z)\\ L_{m_{j}}^{\top}(z)&0_{(m_{j}+1)\times(m_{j}+1)}\end{bmatrix},L_{m_{j}}=[I_{m_{j}}\mathbf{0}]+z[\mathbf{0}\,I_{m_{j}}]\in\mathbb{C}^{m_{j}\times(m_{j}+1)}[z],j\in[k],$
    where $H_{k+1}=\emptyset$ or $H_{k+1}$ is a nonsingular Hermitian pencil. In particular, $\det(\lambda I_{n}-H(z))=\lambda^{k}p(\lambda,z)$, and $p(0,z)\not\equiv 0$. Hence, $H(z)$ has exactly $k$ 0-eigenvalues, and nkn-k eigenvalues that vanish at most nkn-k points in $\mathbb{C}$.
Proof.
(1) is straightforward.
(2) Assume that $H(z)$ is $1$-indecomposable singular Hermitian pencil. Then a variation of the Kronecker canonical form for general pencils gives decomposition (27) [14, Theorem 1] or [11, Theorem 6.1]. Next observe that for $H_{j}(z)$ has exactly on eigenvector corresponding 0 eigenvalue. As $H_{j}(z)$ is a real symmetric matrix it for a real $z$ it follows that 0-eigenvalue is of multiplicity one of $H_{j}(z)$. Hence, 0 is an eigenvalue of multiplicity $k$ in $\oplus_{j=1}^{k+1}H_{j}(z)$. Therefore, $\det(\lambda I_{n}-H(z))=\lambda^{k}p(\lambda,z)$ where the joint degree of $p(\lambda,z)$ is nkn-k. Furhtermore $p(0,z)$ is a nonzero polynomial of degree at most nkn-k. (if $p(0,z)$ vanish identically, we will deduce that 0 is an eigenvalue of multiplicity at $k+1$.) Hence the product of the nonvanishing nkn-k eigenvalues of $H(z)$ vanish at the zeros of $p(0,z)=0$, which is a set of cardinality at most nkn-k. ∎
Example 1.
Let $H(z)=\begin{bmatrix}0&1&z\\ 1&0&0\\ z&0&0\end{bmatrix}$. Then
(28) $\det(\lambda I_{3}-H(z))=\lambda^{3}-(1+z^{2})\lambda,$
$\lambda_{1}(z)=\sqrt{1+z^{2}},\mathbf{u}_{1}=\frac{1}{\sqrt{2(z^{2}+1)}}(\sqrt{z^{2}+1},1,z)^{\top},$
$\lambda_{2}(z)=0,\mathbf{u}_{2}(z)=\frac{1}{\sqrt{z^{2}+1}}(0,-z,1)^{\top},z),$
$\lambda_{3}(z)=-\sqrt{1+z^{2}},\mathbf{u}_{3}=\frac{1}{\sqrt{2(z^{2}+1)}}(-\sqrt{z^{2}+1},1,z)^{\top}.$
The points where the two nonzero eigenvalues are zero are $\pm\sqrt{-1}$. Let $\Omega$ be a simply connected domain in $\mathbb{C}$ containing $\mathbb{R}$ and does not contain $\pm\sqrt{-1}$. Then the eigenvalues and eignevectors are analytic in $\Omega$. The eigenvectors are orthonormal on $\mathbb{R}$. Furthermore,
$\{H(x)>0\}=\mathbf{u}_{1}(x)\mathbf{u}_{1}(x)^{\top},H(x)_{+}=\lambda_{1}(x)\mathbf{u}_{1}(x)\mathbf{u}_{1}(x)^{\top},$
$\{-H(x)>0\}=\mathbf{u}_{3}(x)\mathbf{u}_{3}(x)^{\top},H(x)_{1}=\lambda_{3}(x)\mathbf{u}_{3}(x)\mathbf{u}_{3}(x)^{\top}.$
The eigenvalues and eigenvectors, except $\lambda_{2}(z)$, are two multivalued on $\mathbb{C}\setminus\{\pm\sqrt{-1}\}$.
In what follows we need a following variation of Rellich’s theorem [6, Theorem 4.18.2] :
Theorem 4.
Let $H(z)$ be a Hermitian pencil given by (25). There exists a simply connected domain $\Omega\subset\mathbb{C}$ such that $\mathbb{R}\subset\Omega,\bar{\Omega}=\Omega$, and the following conditions hold: The polynomial $\det(\lambda I_{n}-H(z))$ splits to $\prod_{j=1}^{n}(\lambda-\alpha_{j}(z))$, where each $\alpha_{j}(z)$ is analytic in $\Omega$ and satisfies $\alpha_{j}(\bar{z})=\overline{\alpha_{j}(z)}$ for $z\in\Omega$ and $j\in[n]$. If $\alpha_{j}(z)$ is not identically zero then each has a finite number of zeros in $\Omega$. To each eigenvalue $\alpha_{j}(z)$ corresponds an analytic eigenvector $\mathbf{u}_{j}(z)$: $A(z)\mathbf{u}_{j}(z)=\alpha_{j}(z)\mathbf{u}_{j}(z)$. Furthermore $\langle u_{i}(\bar{z}),u_{j}(z)\rangle=\delta_{ij}$ for $i,j\in[n]$ for $z\in\Omega$.
There exists a finite number of point $Z=\{\zeta_{1},\ldots,\zeta_{m}\}\subset\mathbb{C}\setminus\mathbb{R}$ such that each $\alpha_{j}(z)$ and $\mathbf{u}_{j}(z)$ is a finite multi-valued function on $\mathbb{C}\setminus Z$.
Proof.
All the claims of the theorem are stated and proved in [6, Theorem 4.18.12], except the statement that each $\alpha_{j}(z)$ that is not identically zero has a finite number of zeros. This statement follows from Lemma 2. ∎
Corollary 1.
Let $H(z)$ $n\times n$ nonzero Hermitian pencil. Assume that $Z=\{\zeta_{1},\ldots,\zeta_{m}\},\subset\mathbb{R},0\leqslant m\leqslant n$ be all the real points where the nonzero eigenvalues of $H(x)$ are zero.
  1. (1)
    The projections $\{H(x)>0\},\{-H(x)>0\}$ and $H(x)_{+},H(x)_{-}$ are analytic on the intervals $\mathbb{R}\setminus Z$.
  2. (2)
    The matrices $H_{+}(x),H_{-}(x)$ are continuous on $\mathbb{R}$.
A.2. An infinite dimensional case
Assume that $\mathcal{H}$ is a separable infinite dimensional Hilbert space. Denote by $\mathcal{S}^{\prime}_{++}$ the set positive operaotrs with a bounded inverse. That is, $T\in\mathcal{S}^{\prime}_{++}\iff\langle T\mathbf{x},\mathbf{x}\rangle\geqslant\varepsilon(T)\|\mathbf{x}\|^{2}$ for some $\varepsilon(T)>0$.
Let $\{T_{n},n\in\mathbb{N}\}\subset\mathcal{B}$. On $\mathcal{B}$ we have three topologies:
  1. (1)
    Norm topology: $\lim T_{n}\to T\iff\lim_{n\to\infty}\|T_{n}-T\|=0$.
  2. (2)
    Strong topology: $\lim T_{n}\stackrel{{\scriptstyle strong}}{{\to}}T\iff\lim_{n\to\infty}\|T_{n}\mathbf{x}-T\mathbf{x}\|=0$ for each $\mathbf{x}\in\mathcal{H}$.
  3. (3)
    Weak topology: $\lim T_{n}\stackrel{{\scriptstyle weak}}{{\to}}T\iff\lim_{n\to\infty}\langle\mathbf{y},T_{n}\mathbf{x}\rangle=\langle\mathbf{y},T\mathbf{x}\rangle$ for each $\mathbf{x},\mathbf{y}\in\mathcal{H}$.
Recall that if the sequences $\langle\mathbf{y},T_{n}\mathbf{x}\rangle$ ($\{T_{n}\mathbf{x}\}$) converge for each $\mathbf{x},\mathbf{y}\in\mathcal{H}$ ($\mathbf{x}\in\mathcal{H}$) then $\sup_{n\in\mathbb{N}}\|T_{n}\|<\infty$. Furthermore, there exists $T\in\mathcal{B}$ such that $\lim T_{n}\stackrel{{\scriptstyle weak}}{{\to}}T$ ($\lim T_{n}\stackrel{{\scriptstyle strong}}{{\to}}T$) [13, Sec. VI.1].
Let $T\in\mathcal{S}$. One associates with $T$ a family of projective-valued measures on the interval $\|T\|,\|T\|]$, see [13, VIII.3,Proposition]. Then, for any bounded Borel function ffon $\|T\|,\|T\|]$ one can define $f(T)=\int_{-\|T\|}^{\|T\|}f(\lambda)dP_{\lambda}$ such that $\langle\mathbf{x},f(T)\mathbf{x}\rangle=\int_{-\|T\|}^{\|T\|}f(\lambda)d\langle\mathbf{x},P_{\lambda}\mathbf{x}\rangle$. In particular
(29) $T=\int_{-\|T\|}^{\|T\|}\lambda dP_{\lambda},\quad T_{+}=\int_{0}^{\|T\|}\lambda dP_{\lambda},\quad T_{-}=-\int^{0}_{-\|T\|}\lambda dP_{\lambda},$
$\{T>0\}=\int_{(0,\|T\|]}dP_{\lambda},\quad\{-T>0\}\int_{[-\|T\|,0)}dP_{\lambda},$
$\log T=\int_{\varepsilon(T)}^{\|T\|}\log\lambda dP_{\lambda},\quad\langle\mathbf{x},T\mathbf{x}\rangle\geqslant\varepsilon(T)\|\mathbf{x}\|^{2},\mathbf{x}\in\mathcal{H},\varepsilon(T)>0.$
However, one prefers integral formulas which do not involve projections. We start with the following well known integral formula of $\log b,b>0$: $\log b=\int_{0}^{\infty}\big((1+x)^{-1}-(b+x)^{-1}\big)dx$ which reduces to:
$\log B=\int_{0}^{\infty}\big((Id+x\mathbf{1})^{-1}-(B+x\mathbf{1})^{-1})dx,\quad B\in\mathcal{S}^{\prime}_{++}.$
(Here $\mathbf{1}$ is the identity operator.) Observe that the second integral converge as $\|(B+x\mathbf{1})^{-1}\|\leqslant(\varepsilon(B)+x)^{-1}$ for $B\in\mathcal{S}^{\prime}_{++}(\mathcal{B})$ and $x\geqslant-\varepsilon(B)/2$. Hence,
$\frac{1}{t}\big(\log(B+tA)-\log B\big)=\frac{1}{t}\int_{0}^{\infty}\big((B+x\mathbf{1})^{-1}-(B+tA+x\mathbf{1})^{-1})dx=$
$\frac{1}{t}\int_{0}^{\infty}(B+x\mathbf{1})^{-1}\big((B+tA+x\mathbf{1})-(B+xId)\big)(B+tA+x\mathbf{1})^{-1}dx=$
$\int_{0}^{\infty}(B+x\mathbf{1})^{-1}A(B+tA+x\mathbf{1})^{-1}dx,\quad A\in\mathcal{S},|t|\|A\|\leqslant\varepsilon(B)/2.$
Thus, we deduce
(30) $D[B](A)=\lim_{t\to 0}\frac{1}{t}\big(\log(B+tA)-\log B\big)=\int_{0}^{\infty}(B+x\mathbf{1})^{-1}A(B+x\mathbf{1})^{-1}dx,A\in\mathcal{S},B\in\mathcal{S}^{\prime}_{++}.$
Next recall the formula for $\sqrt{a},a\geqslant 0$: $\sqrt{a}=\frac{2}{\pi}\int_{0}^{\infty}\frac{a}{a+x^{2}}dx$. we deduce
(31) $|A|=\sqrt{A^{2}}=\frac{2}{\pi}\int_{0}^{\infty}{A^{2}}(A^{2}+x^{2}\mathbf{1})^{-1}dx=\frac{2}{\pi}\int_{0}^{\infty}(A^{2}+x^{2}\mathbf{1})^{-1}A^{2}dx,$
$A_{+}=\frac{1}{2}\big(A+\frac{2}{\pi}\int_{0}^{\infty}{A^{2}}(A^{2}+x^{2}\mathbf{1})^{-1}dx\big)$
$A_{-}=\frac{1}{2}\big(-A+\frac{2}{\pi}\int_{0}^{\infty}{A^{2}}(A^{2}+x^{2}\mathbf{1})^{-1}dx\big).$
Next observe that since $\|A^{2}(A^{2}+x^{2}\mathbf{1})^{-1}\|\leqslant\frac{\|A^{2}\|}{\|A^{2}\|+x^{2}}$ the above integral converges.
It was shown by Kato [8], basically using the first equality in (30):
(32) $\||A|-|B|\|\leqslant\frac{2}{\pi}\|A-B\|\big(2+\log\frac{\|A\|+\|B\|}{\|A-B\|}\big),\quad A,B\in\mathcal{B}.$
For $A,B\in\mathrm{S}_{2}$ Araki [1, Lemma 5.2] showed
(33) $\||A|-|B|\|_{2}\leqslant\|A-B\|_{2}.$
In view of (18) we deduce:
(34) $\max(\|A_{+}-B_{+}\|,\|A_{-}-B_{-}\|)\leqslant\frac{1}{\pi}\|A-B\|\big(\frac{\pi+4}{2}+\log\frac{\|A\|+\|B\|}{\|A-B\|}\big),\quad A,B\in\mathcal{B}.$
A.3. Approximation lemmas
Lemma 3.
[…]
$T_{n}=P_{n}TP_{n},\quad T_{n,+}=(T_{n})_{+},\quad T_{n,-}=(T_{n})_{-}\in\mathbb{N}.$
  1. (1)
    Assume that $T\in\mathcal{S}_{\infty}$.
    (35) $\|T_{n,+}\|=\max_{\mathbf{x}\in\mathbf{V}_{n},\|\mathbf{x}\|\leqslant 1}\langle\mathbf{x},T\mathbf{x}\rangle\leqslant\|T_{n+1,+}\|\leqslant\|T_{+}\|,\quad n\in\mathbb{N},$
    $\lim_{n\to\infty}\|T-T_{n}\|=0,\lim_{n\to\infty}\|T_{+}-T_{n,+}\|=0,\quad\lim_{n\to\infty}\|T_{-}-T_{n,-}\|=0.$
  2. (2)
    Assume that $p\in[1,\infty)$ and $T\in\mathcal{S}_{p}$. Then
    (36) $\|T_{n,+}\|_{p}\leqslant\|T_{n+1,+}\|_{p}\leqslant\|T_{+}\|_{p},\quad\|T_{n,-}\|_{p}\leqslant\|T_{n+1,-}\|_{p}\leqslant\|T_{-}\|_{p},\quad n\in\mathbb{N},$
    $\lim_{n\to\infty}\|T-T_{n}\|_{p}=0,\lim_{n\to\infty}\|T_{+}-T_{n,+}\|_{p}=0,\quad\lim_{n\to\infty}\|T_{-}-T_{n,-}\|_{p}=0.$
  3. (3)
    Assume that $T\in\mathcal{S}$. Then the following sequences converge in the strong topology
    (37) $\lim_{n\to\infty}T_{n}=T,\quad\lim_{n\to\infty}T_{n,+}=T_{+},\quad\lim_{n\to\infty}T_{n,-}=T_{-}.$
Proof.
(1) Let $T\in\mathcal{S}_{\infty}$. Then $T=T_{+}-T_{-}$, where $T_{+}\mathcal{H}=\mathcal{H}_{+},T_{-}\mathcal{H}=\mathcal{H}_{-}$, $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$ are two orthogonal subspaces in $\mathcal{H}$. We first consider the generic case where $\mathcal{H}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$, and $\mathcal{H}_{+},\mathcal{H}_{-}$ are infinite dimensional. Assume that $\mathcal{H}_{+}$ and $\mathcal{H}_{-}$ have orthonormal bases $\mathbf{x}_{i},i\in\mathbb{N}$ and $\mathbf{y}_{i},i\in\mathbb{N}$ respectively. Hence,
(38) $T_{+}=\sum_{i=1}^{\infty}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*},\mu_{1}\geqslant\cdots\geqslant\mu_{n}>0,\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle=\delta_{ij},i,j,n\in\mathbb{N},\lim_{n\to\infty}\mu_{n}=0,$
$T_{-}=\sum_{i=1}^{\infty}\nu_{i}\mathbf{y}_{i}\mathbf{y}_{i}^{*},\nu_{1}\geqslant\cdots\geqslant\nu_{n}>0,\langle\mathbf{y}_{i},\mathbf{y}_{j}\rangle=\delta_{ij},i,j,n\in\mathbb{N},\lim_{n\to\infty}\nu_{n}=0$
Observe that the above expansions converge in norm. This follows from the observation:
(39) $\|\sum_{i=j}^{k}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*}\|=\mu_{j}\textrm{ for }k\geqslant j.$
Next observe that the first equality in (35) is obvious. (Note that $T_{n,+}$ can be 0.) Hence, $\|T_{n,+}\|\leqslant\|T_{n+1,+}\|$. Next observe that $T_{n,+}=Q_{n}T_{n}Q_{n}$, where $Q_{n}:\mathbf{V}_{n}\to\mathbf{V}_{n}$ is a projection. Thus
(40) $T_{n,+}=Q_{n}(P_{n}T_{+}P_{n})Q_{n}-Q_{n}(P_{n}T_{-}P_{n})Q_{n}\leqslant Q_{n}(P_{n}T_{+}P_{n})Q_{n}\Rightarrow\|T_{n,+}\|\leqslant\|T_{+}\|.$
We now prove the equality $\lim_{n\to\infty}\|T_{+}-P_{n}T_{+}P_{n}\|=0$. Define
(41) $R_{N}=\sum_{i=N+1}^{\infty}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*}\in\mathcal{S}_{\infty,+},\quad N\in\mathbb{N}.$
Fix $\delta>0$. Choose $N(\delta)$ such that $\mu_{N}(\delta)<\delta/4$. Then
(42) $\|R_{N(\delta}\|<\delta/4\Rightarrow\|P_{n}R_{N(\delta)}P_{n}\|<\delta/4\Rightarrow$
$\|T_{+}-P_{n}T_{+}P_{n}\|<\|\sum_{i=1}^{N(\delta)}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*}-P_{n}\big(\sum_{i=1}^{N(\delta)}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*}\big)P_{n}\|+\delta/2.$
As $P_{n}(\mathbf{x}_{i}\mathbf{x}_{i}^{*})P_{n}=(P_{n}\mathbf{x}_{i})(P_{n}\mathbf{x}_{i})^{*}\to\mathbf{x}_{i}\mathbf{x}_{i}^{*}$ in norm we deduce that there exists $M(\delta)$ such that for $n>M(\delta)$ we have $\|T_{+}-P_{n}T_{+}P_{n}\|<\delta$. Similarly, $\lim_{n\to\infty}\|T_{-}-P_{n}T_{-}P_{n}\|=0$. This show that $\lim_{n\to\infty}\|T-T_{n}\|=0$. Kato’s inequality (34) yields the last two inequalities in (35). The other cases are proved similarly
(2) Assume that $p\in[1,\infty)$. Assume first the generic case as in part (1). Then $\|T\|_{p}=(\|T_{+}\|^{p}_{p}+\|T_{-}\|^{p}_{p})^{1/p}$. The inequality $\|T_{n,+}\|_{p}\leqslant\|T_{n+1,+}\|_{p}$ follows from the Cauchy interlacing theorem. Indeed, we can assume that $T_{n}\in\mathrm{H}_{n},T_{n+1}\in\mathrm{H}_{n+1}$, and $T_{n}$ is obtained from $T_{n+1}$ by deleting the last row and column. Assume that $\lambda_{1}(T_{n+1})\geqslant\ldots\geqslant\lambda_{m}(T_{n+1})\geqslant 0>\lambda_{m+1}(T_{n+1})$, where $m\leqslant n+1$. Cauchy interlacing theorem yields:
(43) $\lambda_{i}(T_{n+1})\geqslant\lambda_{i}(T_{n})\geqslant\lambda_{i-1}(T_{n+1}),\quad i\in[n].$
Hence, $\lambda_{m}(T_{n})\in[\lambda_{m}(T_{n+1}),\lambda_{m+1}(T)]$. Suppose first that $\lambda_{m}(T_{n})\in[\lambda_{m}(T_{n+1},0]$. Then
$\lambda_{i}(T_{n+1,+})=\lambda_{i}(T_{n+1}),i\in[m],\quad\lambda_{i}(T_{n+1,+})=0\textrm{ for }i>m,$
$\lambda_{i}(T_{n,+})=\lambda_{i}(T_{n}),i\in[m],\quad\lambda_{i}(T_{n,+})=0\textrm{ for }i\geqslant m+1\Rightarrow$
$\|T_{n,+}\|_{p}=\big(\sum_{i=1}^{m-1}\lambda_{i}^{p}(T_{n})\big)^{1/p}\leqslant\|T_{n+1,+}\|_{p}=\big(\sum_{i=1}^{m}\lambda_{i}^{p}(T_{n+1})\big)^{1/p}$
Similar arguments yield that $\|T_{n,-}\|_{p}\leqslant\|T_{n+1,-}\|_{p}$. The above inequalities hold when $\lambda_{m}(T)\in(0,\lambda_{m+1}(T_{n+1})]$.
Our assumption that $T_{+}$ has infinite number of eigenvalues and the fact that $\lim_{n\to\infty}\|T_{+}-T_{n,+}\|=0$ yields the following fact. Fix $k\geqslant 1$ and consider the sequence $\lambda_{i}(T_{n}),n\in\mathbb{N}$ for $i\in[k]$. Then each $\lambda_{i}(T_{n}),n\geqslant k$ is a nondecreasing sequence that converges to $\mu_{i}$ for $i\in[k]$. In particular, we deduce that
(44) $\sum_{i=j}^{n}\lambda_{i}^{p}(T_{n,+})\leqslant\sum_{i=j}^{n}\mu_{i}^{p}(T_{+}),\quad j\in[n],n\in\mathbb{N}\Rightarrow$
$\|T_{n,+}\|_{p}\leqslant\big(\sum_{i=j}^{n}\mu_{i}^{p}(T_{+})\big)^{1/p}\leqslant\|T_{+}\|_{p}.$
Similar results hold for $\|T_{n,-}\|_{p}$.
We now show that $\lim_{n\to\infty}\|T_{+}-T_{n,+}\|_{p}=0$. Fix $\delta>0$. Let $R_{N}$ be defined by (41). Clearly, $R_{N}\in\mathcal{S}_{p,+}$ as as $\|R_{N}\|_{p}=\big(\sum_{i=N+1}^{\infty}\mu_{i}^{p}\big)^{1/p}$. Choose $N(\delta)$ such that $\|R_{N}\|_{p}<\delta/4$. Clearly $P_{n}R_{N}P_{n}\geqslant 0$. Hence, the first set of inequalities of (36) yield $\|P_{n}R_{N}P_{n}\|_{p}\leqslant\|R_{N}\|_{p}<\delta/4$. As in (42) we obtain
$\|T_{+}-P_{n}T_{+}P_{n}\|_{p}<\|\sum_{i=1}^{N(\delta)}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*}-P_{n}\big(\sum_{i=1}^{N(\delta)}\mu_{i}\mathbf{x}_{i}\mathbf{x}_{i}^{*}\big)P_{n}\|_{p}+\delta/2.$
It is left to show rank-two operator $\mathbf{x}_{i}^{*}\mathbf{x}_{i}-(P_{n}\mathbf{x}_{i})(P_{n}\mathbf{x}_{i})^{*}$ converges in the operator norm $\|\cdot\|_{p}$ to zero, which is trivial. Hence, $\|T_{+}-T_{n,+}\|_{p}<\delta$ for $n>M(\delta)$. This proves the last two equalities in (35), which imply that $\lim_{n\to\infty}\|T-T_{n}\|_{p}=0$. Other cases are proved similarly.
(3) Clearly, $P_{n}\stackrel{{\scriptstyle strong}}{{\to}}\mathbf{1}$. Recall [13, Sec. VI, Problem (6d)] that if $A_{n},B_{n}\in\mathcal{S}$ converge in the strong topology to $A,B\in\mathrm{S}(\mathcal{B})$ respectively, then $A_{n}B_{n}\stackrel{{\scriptstyle strong}}{{\to}}AB$. Hence
$T_{n,+}-T_{n,-}=P_{n}TP_{n}\stackrel{{\scriptstyle strong}}{{\to}}T,=T_{+}-T_{-}\quad,(P_{n}TP_{n})^{2}\stackrel{{\scriptstyle strong}}{{\to}}T^{2}.$
Problem 14(b) in [13, Sec. VI] states that if $A_{n}\in\mathrm{S}_{+}(\mathcal{B})$ converges strongly to $A\in\mathrm{S}_{+}(\mathcal{B})$ then $\sqrt{A_{n}}\stackrel{{\scriptstyle strong}}{{\to}}\sqrt{A}$. Hence
$T_{n,+}+T_{n,-}=\sqrt{T_{n}^{2}}\stackrel{{\scriptstyle strong}}{{\to}}\sqrt{T^{2}}=T_{+}+T_{-}.$
These two equalities prove the last two equalities in (37). ∎